OpenStudy (anonymous):

PM and PM are tangent to a circle O at M and N, respectively; m< MON=120 degrees, and OM= ON=5. Find the perimeter of the region PMN makes.

OpenStudy (anonymous):

Answer choices are: 10+10 pi, \[5\sqrt{3} +10 \pi\] \[5\sqrt{3} +10\pi/3 ,\] \[10\sqrt{3} +10\pi/3 \] or \[5\sqrt{3} +10pi\]